# Heterochromatic Geometric Transversals of Convex sets

### v4

@@ 3 Heterochromatic $(\aleph_{0},\cdot)$-theorem for fat convex sets
Observe that Lemma 3.4 is equivalent to the following result:
Lemma 3.5 (Contraposition of Lemma 3.4).
For each $n\in\mathbb{N}$, suppose that $\mathcal{F}_{n}$ is a family of closed unit balls in $\mathbb{R}^{d}$ and the collection $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies the weak heterochromatic $(\aleph_{0},2)$-property. Then there exists an $i\in\mathbb{N}$ such that the family $\mathcal{F}_{i}$ can be pierced by finitely many points.
Now applying Lemma 3.1, we obtain the following:
Corollary 3.6 (Heterochromatic $(\aleph_{0},2)$-theorem for closed unit balls).
For each $n\in\mathbb{N}$, suppose that $\mathcal{F}_{n}$ is a family of units balls in $\mathbb{R}^{d}$ and the collection $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies the heterochromatic $(\aleph_{0},2)$-property. Then all but finitely many families are pierceable by finitely many points.
Now we are ready to prove Theorem 1.10 $(ii)$.
Theorem 3.7 (Restatement of Theorem 1.10 $(ii)$).
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of units balls in $\mathbb{R}^{d}$ such that $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ satisfy the weak heterochromatic $(\aleph_{0},2)$ property. Then there are at least two families $\mathcal{F}_{i},\mathcal{F}_{j},\;i\neq j$, each of which is pierceable by finite number of points.
@@
Next, we shall show the tightness of the above result, that is Proposition 1.11. See 1.11
@@
Suppose $\mathcal{F}_{1}$ is the family containing only one unit ball $B_{1}$ in $\mathbb{R}^{2}$ with center at the origin and $\mathcal{F}_{2}$ contains the $4$ unit balls $B_{2,1},\dots,B_{2,4}$ in $\mathbb{R}^{2}$ with centers at the points $(1.5,0)$, $(-1.5,0)$, $(0,1.5)$, and $(0,-1.5)$ in $\mathbb{R}^{2}$ respectively. Then $\forall B\in\mathcal{F}_{2},\;B\cap B_{1}\neq\emptyset$ and clearly each of the families $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ are pierceable by a finite number of points. Now for each $n\in\mathbb{N},\;n\geq 3$, we define $\mathcal{F}_{n}$ to be the family of unit balls $\{B_{n,j}\;|\;j\in\mathbb{N}\}$, where $B_{n,j}$ is the unit ball in $\mathbb{R}^{d}$ with center at $c_{n,j}=(4j,4n),\;\forall j\in\mathbb{N}$. Then for any $m,n,i,j\in\mathbb{N},\;m,n\geq 3$ with $(m,i)\neq(n,j)$ we have $B_{m,i}\cap B_{n,j}=\emptyset$ and for each $n\in\mathbb{N},\;n\geq 3$, the family $\mathcal{F}_{n}$ is not pierceable by finite number of points. Hence $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ satisfies both the properties required.
Finally, we shall prove Theorem 1.10 $(i)$ for closed unit balls. Lemma 3.1 ensures that it is sufficient to prove the following:
Theorem 3.8.
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of closed unit balls in $\mathbb{R}^{d}$ such that $\mathcal{F}_{n}$ is not pierceable by a finite number of $k$-flats, where $10\leq k\leq d-$ 1. Then there exists a $k$-independent heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of closed unit balls.
Now to prove Theorem 3.8 we will need the following lemmas.
Lemma 3.9.
Suppose $\mathcal{F}$ is a family of closed balls in $\mathbb{R}^{d}$ such that $\mathcal{F}$ is not pierceable by a finite number of $k$-flats. Then there exists $\hat{x}\in\mathbb{S}^{d-1}$ such that $\forall\varepsilon>0,\mathcal{F}_{(\hat{x},\varepsilon)}$ is not pierceable by finite number of $k$-flats.
@@
The proof of this is identical to Keller and Perles’[KP22] proof for unit balls. Regardless, we add the proof here for the reader’s convenience. If possible let the statement of the lemma be false. Then $\forall\hat{y}\in\mathbb{S}^{d-1},\exists\varepsilon(\hat{y})$ for which $\mathcal{F}_{(\hat{y},\varepsilon(\hat{y}))}$ is pierceable by finite number of $k$-flats. Now $\left\{B(\hat{y},\varepsilon(\hat{y}))\;|\;\hat{y}\in\mathbb{S}^{d-1}\right\}$ is an open cover of $\mathbb{S}^{d-1}$. So, by compactness of $\mathbb{S}^{d-1},\;\exists\hat{y}_{1},\dots,\hat{y}_{r}\in\mathbb{S}^{d-1}$ such that $\mathbb{S}^{d-1}\subseteq\bigcup_{i=1}^{r}B(\hat{y}_{i},\varepsilon(\hat{y}_{i}))$ and hence $\mathcal{F}\subseteq\bigcup_{i=1}^{r}\mathcal{F}_{(\hat{y}_{i},\varepsilon(\hat{y}_{i}))}$. For each $i\in\ [r]$, since $\mathcal{F}_{(\hat{y}_{i},\varepsilon(\hat{y}_{i}))}$ is pierceable by finitely many $k$-flats, we get $\mathcal{F}$ is pierceable by finitely many $k$-flats, a contradiction. Hence the result. ∎
The following important lemma is due to Keller and Perles.
Lemma 3.10 ([KP22]).
Let $\mathcal{F}$ be a family of closed unit balls in $\mathbb{R}^{d}$, $10\leq k\leq d-$ 1 and $m\in\mathbb{N}$. If every finite subset of $\mathcal{F}$ can be pierced by $m$ $k$-flats, then $\mathcal{F}$ can be pierced by $m$ $k$-flats.
Lemma 3.11.
[…]
Proof.
Consider the space $C=B_{1}\times B_{2}\times\dots\times B_{k+1}$. Then clearly $C$ is compact. Now take any $c=(b_{1}\dots,b_{k+1})\in C$ and suppose $\kappa_{c}$ is the $k$-flat generated by $\{b_{1}\dots,b_{k+1}\}$ in $\mathbb{R}^{d}$. We define a function $f\;:\;C\rightarrow\mathbb{R}_{\geq 0}$ such that
[…]
Then $f$ is a continuous function. Since for all $\kappa\in\mathcal{K}$, direction of $\kappa$ does not contain the point $\hat{x}$, we have $f(c)>0$, for all $c\in C$. So by compactness of $C$ we get that, there exists $\varepsilon>0$ such that $f(c)\geq\varepsilon$, for all $c\in C$. ∎
Now we are ready to prove Theorem 3.8.
See 3.8
Proof.
For each $n\in\mathbb{N}$, since $\mathcal{F}_{n}$ is not pierceable by finitely many $k$-flats, by Lemma 3.9, there exists $\hat{x}_{n}\in\mathbb{S}^{d-1}$ such that $\forall\varepsilon>0,\mathcal{F}_{n(\hat{x}_{n},\varepsilon)}$ is not pierceable by finite number of $k$-flats.
Now by compactness of $\mathbb{S}^{d-1}$, $\exists\hat{x}\in\mathbb{S}^{d-1}$ such that $\{\hat{x}_{n}\}_{n\in\mathbb{N}}$ has a subsequence converging to $\hat{x}$. Without loss of generality, we may assume that the convergent subsequence itself is $\{\hat{x}_{n}\}$. Then for every $\varepsilon>0$ there exists $N(\varepsilon)\in\mathbb{N}$ such that $\forall n\geq N(\varepsilon),\;\text{dist}(\hat{x},\hat{x}_{n})<\varepsilon$.
Now we shall induct on $k$ and $d$ to prove the result. If $k=0$, then by Lemma 3.4 we get the result. So we let $k>0$. Suppose that the result holds for families of closed unit balls in $\mathbb{R}^{d-1}$ with respect to $(k-1)$-flats. We shall show that it holds for $k$-flats in $\mathbb{R}^{d}$.
Choose a sequence $\{\varepsilon_{n}\}_{n\in\mathbb{N}}$ of strictly decreasing positive numbers such that $\varepsilon_{n}\to 0$ as $n\to\infty$.
Now we show that we can pick an infinite sequence $\{B^{\prime}_{n}\}_{n\in\mathbb{N}}$ of unit balls such that $\{B^{\prime}_{n}\}_{n\in\mathbb{N}}$ cannot be pierced by finitely many $k$-flats, and for every $i\in\mathbb{N}$, there is an $m(i)\in\mathbb{N}$ for which we have $\bigcup_{n=i}^{\infty}\mathcal{F}_{n(\hat{x}_{n},\varepsilon_{n})}\supsetneq\{B^{\prime}_{n}\}_{n=m(i)}^{\infty}$ with the property that for all $n\geq m(i)$, $B^{\prime}_{n}\cap B(\mathcal{O},i)=\emptyset$. For all $n\in\mathbb{N}$, let $\mathcal{F}^{\prime}_{n(\hat{x}_{n},\varepsilon_{n})}=\{B\in\mathcal{F}_{n(\hat{x}_{n},\varepsilon_{n})}\mid B\cap B(\mathcal{O},n)=\emptyset\}$. Then clearly, $\mathcal{F}^{\prime}_{n(\hat{x}_{n},\varepsilon_{n})}$ cannot be pierced by finitely many $k$-flats. Let us have $\mathcal{B}_{1},\dots,\mathcal{B}_{i-1}$ such that $\forall j\in[i-1]$, we have $\mathcal{B}_{j}\subsetneq\mathbb{F}^{\prime}_{j(\hat{x}_{j},\varepsilon_{j})}$ and $\mathcal{B}_{j}$ is finite. Then there is an $M(i-1)\in\mathbb{N}$ such that at least $M(i-1)$ $k$-flats are needed to pierce every set in $\bigcup_{j=1}^{i-1}\mathcal{B}_{j}$. We show that there is a finite set $\mathcal{B}_{i}\subsetneq\mathcal{F}^{\prime}_{i(\hat{x}_{i},\varepsilon_{i})}$ such that at least $M(i)\geq M(i-1)+1$ $k$-flats are needed to pierce every set in $\bigcup_{j=1}^{i}\mathcal{B}_{j}$. If not, then every finite subset of the set $\mathcal{B}_{1}\cup\dots\cup\mathcal{B}_{i-1}\cup\mathcal{F}^{\prime}_{i(\hat{x}_{i},\varepsilon_{i})}$, and therefore $\mathcal{F}^{\prime}_{i(\hat{x}_{i},\varepsilon_{i})}$ has a $k$-transversal of size $M(i-1)$. Then by Lemma 3.10, $\mathcal{F}^{\prime}_{i(\hat{x}_{i},\varepsilon_{i})}$ is pierceable by $m(i-1)$ $k$-flats, which is a contradiction. Therefore, we get a finite set $\mathcal{B}_{i}\subsetneq\mathcal{F}^{\prime}_{i(\hat{x}_{i},\varepsilon_{i})}$ for which $\bigcup_{j=1}^{i}\mathcal{B}_{j}$ cannot be pierced by fewer than $M(i)$ $k$-flats. Continuing this way, we get our sequence $\{B^{\prime}_{n}\}_{n\in\mathbb{N}}=\bigcup_{j=1}^{\infty}\mathcal{B}_{j}$ which is not pierceable by finitely many $k$-flats. Let =$\ell=\{R\hat{x}\mid R\geq 0\}$. Let $S$ denote the $(d-1)$-flat through $\mathcal{O}$ and perpendicular to $\ell$, and let for every $B^{\prime}_{n}$, $\pi(B^{\prime}_{n})$ denote the projection of $B^{\prime}_{n}$ onto $S$. The sequence $\{\pi(B^{\prime}_{n})\}_{n\in\mathbb{N}}$ then does not have a finite $(k-1)$-transversal, because then we would have finitely many $k$-flats parallel to $\ell$ piercing $\{B^{\prime}_{n}\}_{n\in\mathbb{N}}$. So, there is a subsequence of $\{B^{\prime}_{n}\}_{n\in\mathbb{N}}$ whose projection onto $S$ is $(k-1)$-independent. Without loss of generality, let that subsequence be $\{B^{\prime}_{n}\}_{n\in\mathbb{N}}$. For any $k+1$ sets $B_{1},\dots,B_{k+1}$ picked from $\{B^{\prime}_{n}\}_{n\in\mathbb{N}}$, there is no $k$-flat parallel to $\ell$ passing through every $B_{i}$, $11\leq i\leq k+$ 1. That is because no $(k-1)$-flat passes through the projections of $B_{1},\dots,B_{k+1}$ on $S$, since the intersection of a $k$-flat perpendicular to $\ell$ with $S$ is a $(k-1)$-flat on $S$. Now let $B_{1},\dots,B_{m}$ be picked from $\{B^{\prime}_{n}\}_{n\in\mathbb{N}}$ with $B_{j}\in\mathcal{B}_{n_{j}}$, $n_{j}\neq n_{j}^{\prime}$ for $j\neq j^{\prime}$, $m\geq k+1$, such that no $k$-flat intersects more than $k+1$ distinct $B_{i}$’s and no $k$-flat perpendicular to $\ell$ passes through $k+1$ distinct $B_{i}$’s. Then $B_{1},\dots,B_{m}$ satisfy the conditions of Lemma 3.11, and therefore, there exists an $\varepsilon>0$ such that the direction of every $k$-flat intersecting all of $B_{1},\dots,B_{m}$ is at least $\varepsilon$-distance apart from $\hat{x}$. There is an $i_{1}\in\mathbb{N}$ such that for all $n\geq i_{1}$, $B(\hat{x}_{n},\varepsilon_{n})\subset B(\hat{x},\frac{\varepsilon}{2})$. Corresponding to this $i_{1}$, there is an $m(i_{1})\in\mathbb{N}$ such that $\bigcup_{n=i_{1}}^{\infty}\mathcal{F}^{\prime}_{n(\hat{x}_{n},\varepsilon_{n})}\supsetneq\{B^{\prime}_{n}\}_{n=m(i_{1})}^{\infty}$. Also, let $i_{2}=\max\{n_{1},\dots,n_{m}\}$, where $B_{j}\in\mathcal{B}_{n_{j}}$. Now, it is easy to check that there is a positive integer $i_{3}>0$ for which no $k$-flat that passes through any $k+1$ distinct $B_{i}$’s chosen from $B_{1},\dots,B_{m}$ intersects any sets in $\mathcal{F}^{\prime}_{n(\hat{x}_{n},\varepsilon_{n})}$ whenever $n>i_{3}$, since the angle that any such $k$-flat makes with $\ell$ is bounded by positive numbers. Set $N=1+\max\{i_{1},i_{2},i_{3}\}$, and set $B_{m+1}$ to be any set from $\{B^{\prime}_{n}\}_{n=m(N)}^{\infty}$. Continuing this way, we obtain our heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}}$.
It is quite intuitive from the proof technique of Theorem 3.8 that for higher order transversals, that is for $k$-transversals with $k>0$, the existence of the strictly heterochromatic sequence of balls is not necessarily ensured if the radius of the balls tends to zero. Actually we can show that even if all the balls have a fixed radius, then also the existence of the strictly heterochromatic sequence of balls is not guaranteed.
See 1.12
Figure 2: Dotted region denotes $A_{\{j_{1},j_{2}\}}$
Proof.
4 Heterochromatic $(\aleph_{0},2)$-Theorem for convex sets
In this section, we will prove Theorem 1.14 and Theorem 1.16. To prove Theorem 1.14, we will begin by showing that it is sufficient to establish the following result.
Theorem 4.1.
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of closed convex sets in $\mathbb{R}^{d}$ with $\sigma(\mathcal{F}_{n})=\sigma_{n}$ such that $\sigma_{n}\geq\sigma_{0}>0$ and $\mathcal{F}_{n}$ is not pierceable by finite number of points. Then $\exists$ a heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ that is 00-independent.
Observation 4.2.
Theorem 4.1 implies Theorem 1.14.
Proof.
Suppose Theorem 4.1 holds and for each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of closed convex sets in $\mathbb{R}^{d}$ satisfying heterochromatic $(\aleph_{0},2)$-property and
$\liminf\limits_{n\to\infty}\sigma(\mathcal{F}_{n})>0.$
Then there exists $\sigma_{0}>0$ and a subsequence $\left\{\sigma_{n_{i}}\right\}_{i\in\mathbb{N}}$ such that $\forall i\in\mathbb{N},\;\sigma_{n_{i}}>\sigma_{0}$ and applying Theorem 4.1 on the families $\{\mathcal{F}_{n_{i}}\}_{i\in\mathbb{N}}$, we get that there exists $i\in\mathbb{N}$ such that $\mathcal{F}_{n_{i}}$ is pierceable by finite number of points. Without loss of generality, we may assume that $\mathcal{F}_{n_{1}}$ is pierceable by a finite number of points.
Again applying the same argument over the families $\{\mathcal{F}_{n_{i}}\}_{i>1}$, we get that there exists $i\in\mathbb{N}$ with $i>1$ such that $\mathcal{F}_{n_{i}}$ is pierceable by finite number of points. Continuing this we get that there are infinitely many $i\in\mathbb{N}$ such that $\mathcal{F}_{n_{i}}$ we get that there exists $i\in\mathbb{N}$ such that $\mathcal{F}_{n_{i}}$ is pierceable by a finite number of points. ∎
Now to prove Theorem 4.1 we will need the following technical lemma:
Lemma 4.3.
[…]
Proof.
For each $x\in\mathcal{B}$, suppose $\mathcal{F}_{x}=\{S\in\mathcal{F}\;|\;\text{dist}(x,S)<\delta_{x}\}$. Since for each $S\in\mathcal{F}_{x}$ either $r(S)\geq\delta_{x}\sigma_{0}$ or $x\in S$, $\mathcal{F}_{x}$ is pierceable by finite number of points, say by $t_{x}=t_{x}(\delta_{x})$ points. Now consider the open cover $\{B^{o}(x,\delta_{x})\;|\;x\in\mathcal{B}\}$ of $\mathcal{B}$. Then by the compactness of $\mathcal{B}$, this open cover has a finite subcover, that is, $\exists n\in\mathbb{N}$ such that $\mathcal{B}\subseteq\bigcup_{i=1}^{n}B^{o}(x_{i},\delta_{x_{i}})$. Now we claim that $\mathcal{F}\subseteq\bigcup_{i=1}^{n}\mathcal{F}_{x_{i}}$. Take any $S\in\mathcal{F}$. Since $S\cap\mathcal{B}\neq\emptyset$, $\exists x\in\mathcal{B}\cap S$. Now $\exists i\in[n]$ such that $x\in B^{o}(x_{i},\delta_{x_{i}})$ and hence $\text{dist}(S,x_{i})<\delta_{x_{i}}\implies S\in\mathcal{F}_{x_{i}}$. So $\mathcal{F}\subseteq\bigcup_{i=1}^{n}\mathcal{F}_{x_{i}}$. For each $i\in[n]$, since $\mathcal{F}_{x_{i}}$ is pierceable by $t_{x_{i}}$ points, so we get that $\mathcal{F}$ is pierceable by $\sum_{i=1}^{n}t_{x_{i}}$ points. ∎
Since for each $S\in\mathcal{F}$, $\frac{r(S)}{R(S)}\geq\sigma_{0}$, the following corollary directly from the above lemma.
Corollary 4.4.
[…]
Now we are ready to prove Theorem 4.1. See 4.1
Proof.
For any $m\in\mathbb{N}$, suppose $S_{1},\dots,S_{m}$ are pairwise disjoint sets such that each $S_{i}$ comes from a different family. Without loss of generality, we assume that $\forall i\in[m],\;S_{i}\in\mathcal{F}_{i}$. We shall show that either $\exists S_{m+1}\in\mathcal{F}_{p}$, for some $p>m$, such that $\forall i\in[m],\;S_{i}\cap S_{m+1}=\emptyset$ or we can identify an infinite sequence satisfying all the required conditions.
In the first case, that is, if there exists $S_{m+1}\in\mathcal{F}_{p}$, for some $p>m$, such that $\forall i\in[m],\;S_{i}\cap S_{m+1}=\emptyset$, then there is nothing to prove. Otherwise, $\forall p>m$ and $\forall S\in\mathcal{F}_{p}$, there exists $i\in[m]$ such that $S\cap S_{i}\neq\emptyset$. Then for each $p>m,\;\exists i_{p}\in[m]$ such that infinitely many members of $\mathcal{F}_{p}$ meet $S_{i_{p}}$. So $\exists\ell\in[m]$ and an infinite subset $\mathcal{J}\subseteq\{p\in\mathbb{N}\;|\;p>m\}$ such that for each $j\in\mathcal{J}$, there exists an infinite sequence $\left\{S^{(j)}_{i_{n}}\right\}_{n\in\mathbb{N}}$ satisfying the following:
  • $\forall n\in\mathbb{N},\;S^{(j)}_{i_{n}}\in\mathcal{F}_{j}$ and
  • $\forall j\in\mathcal{J},\;\forall n\in\mathbb{N},\;S^{(j)}_{i_{n}}\cap S_{\ell}\neq\emptyset$.
Without loss of generality we assume that $\forall j\in\mathcal{J},\;\mathcal{F}_{j}=\left\{S^{(j)}_{i_{n}}\right\}_{n\in\mathbb{N}}$ and $\ell=1$.
For each $j\in\mathcal{J}$, since $\mathcal{F}_{j}$ is not pierceable by finite number of points and $\forall S\in\mathcal{F}_{j},\;S\cap S_{1}\neq\emptyset$, so using Corollary 4.4, we get the following: $\forall j\in\mathcal{J},\;\exists x_{j}\in S_{1}$ such that $\forall\delta>0\;\exists S_{j}\in\mathcal{F}_{j}$ satisfying: (a) $\text{dist}(S_{j},x_{j})<\delta$, (b) out-radius of the set $S_{j}$, $R(S_{j})<\delta$ and (c) $x_{j}\not\in S_{j}$.
Then by compactness of $S_{1}$, $\exists x\in S_{1}$ such that $\{x_{j}\}_{j\in\mathcal{J}}$ has a subsequence converging to $x$. Without loss of generality, we assume that the whole sequence $\{x_{j}\}_{j\in\mathcal{J}}$ converges to $x$. We also assume that $\mathcal{J}=\{j_{n}\;|\;n\in\mathbb{N}\}$ and $j_{p}<j_{q}\text{ if and only if }p<q$ and $\forall j\in\mathcal{J},\;x_{j}\neq x$, since the case where $x_{j}=x$ for infinitely many values of $j$ is trivially inferred from Corollary 4.4. We shall now show that there exists an infinite sequence of sets satisfying all the required properties.
Suppose $\Delta_{1}:=\text{dist}(x_{j_{1}},x)\;\mbox{and}\;\delta_{1}:=\frac{\Delta_{1}}{4}.$ Then $\exists S_{j_{1}}\in\mathcal{F}_{j_{1}}$ such that $\text{dist}(S_{j_{1}},x_{j_{1}})<\delta_{1}$, and out-radius of $S_{j_{1}}$, $R(S_{j_{1}})<\delta_{1}$ and $x_{j_{1}}\not\in S_{j_{1}}$. Then we get that, $x\not\in S_{j_{1}}$, that is, $\text{dist}(x,S_{j_{1}})>0$. For otherwise, $x\in S_{j_{1}},\text{ that implies }\text{dist}(c,x)<R(S_{j_{1}})<\delta_{1},$ where $S\subseteq B(c,R(S_{j_{1}}))$. Then
$\Delta_{1}$ $=\text{dist}(x,x_{j_{1}})\leq\text{dist}(x,c)+\text{dist}(c,x_{j_{1}})\leq\text{dist}(x,c)+\text{dist}(S_{j_{1}},x_{j_{1}})+R(S_{j_{1}})<3\delta_{1},$
that is, $\Delta_{1}<3\delta_{1}$, a contradiction.
Since the sequence $\{x_{j}\}_{j\in\mathcal{J}}$ converges to $x$, $\exists j_{p}\in\mathcal{J}$ such that $\text{dist}(x,x_{j_{p}})<\text{dist}(x,S_{j_{1}})/2.$ Without loss of generality, we assume that $p=2$. Suppose $\Delta_{2}:=\text{dist}(x,x_{j_{2}})\;\mbox{and}\;\delta_{2}:=\frac{\Delta_{2}}{4}.$ Then $\exists S_{j_{2}}\in\mathcal{F}_{j_{2}}$ such that $\text{dist}(S_{j_{2}},x_{j_{2}})<\delta_{2}$, out-radius of $S_{j_{2}}$, $R(S_{j_{2}})<\delta_{2}$ and $x_{j_{2}}\not\in S_{j_{2}}$. Then similarly we can show that $x\not\in S_{j_{2}}$, that is, $\text{dist}(x,S_{j_{2}})>0$. Also, we have $S_{j_{1}}\cap S_{j_{2}}=\emptyset$. For otherwise, $\exists z\in S_{j_{1}}\cap S_{j_{2}}$. Then we have,
$2\Delta_{2}=2\text{dist}(x,x_{j_{2}})$ $<\text{dist}(x,S_{j_{1}})$
$=\min\limits_{s\in S_{j_{1}}}\text{dist}(x,s)$
$\leq\text{dist}(x,z)$
$\leq\text{dist}(x,x_{j_{2}})+\text{dist}(x_{j_{2}},z)$
$\leq\text{dist}(x,x_{j_{2}})+\text{dist}(x_{j_{2}},S_{j_{2}})+R(S_{j_{2}})$
$<\Delta_{2}+2\delta_{2}=3\Delta_{2}/2,$
that is, $\Delta_{2}<3\Delta_{2}/4$, a contradiction. Continuing this procedure, we will get an infinite heterochromatic sequence of pairwise disjoint sets. ∎
Observe that if $\mathcal{F}$ is a family of closed balls in $\mathbb{R}^{d}$, then the condition number of $\mathcal{F}$ is $1$. This gives us the following corollary.
Corollary 4.5 (Heterochromatic $(\aleph_{0},2)$-Theorem for closed balls).
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of closed balls in $\mathbb{R}^{d}$ (with no restriction on the radii) and $\mathcal{F}_{n}$ is not pierceable by a finite number of points. Then there exists a heterochromatic sequence $\left\{B_{n}\right\}_{n\in\mathbb{N}}$ of pairwise disjoint balls.
Again the following proposition is a special case of the above theorem.
Proposition 4.6 (Heterochromatic $(\aleph_{0},2)$-theorem for intervals).
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of closed intervals in $\mathbb{R}$ such that $\mathcal{F}_{n}$ is not pierceable by a finite number of points. Then there exists a heterochromatic sequence $\{I_{n}\}_{n\in\mathbb{N}}$ of $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ of pairwise disjoint intervals.
We now prove Theorem 1.16 using the above proposition.
@@ 4 Heterochromatic $(\aleph_{0},2)$-Theorem for convex sets
Proof.
Suppose for any box $B$ and $i\in[d]$, $\pi_{i}(B)$ is the projection of $B$ on the $i$-th axis and $\mathcal{F}_{n,i}=\left\{\pi_{i}(B)\;|\;B\in\mathcal{F}_{n}\right\}$, that is, $\mathcal{F}_{n,i}$ is the collection of closed intervals on real line obtained by projecting the members of $\mathcal{F}_{n}$ on $i$-th axis. Since for each $n\in\mathbb{N}$, $\mathcal{F}_{n}$ is not pierceable by a finite number of points, there exists $i\in[d]$ such that $\mathcal{F}_{n,i}$ is not pierceable by a finite number of points. So there exists an $i\in[d]$ and an infinite subset $\mathcal{I}\subset\mathbb{N}$ such that $\forall n\in\mathcal{I}$, $\mathcal{F}_{n,i}$ is not pierceable by a finite number of points. Then, applying Proposition 4.6, we get that $\exists$ a heterochromatic infinite sequence of pairwise disjoint intervals $\left\{\pi_{i}(B_{n})\right\}_{n\in\mathcal{I}^{\prime}}$ such that $\mathcal{I}^{\prime}\subseteq\mathcal{I}$ is infinite. Then the corresponding boxes $B_{n}\in\mathcal{F}_{n},n\in\mathcal{I}^{\prime}$ form a desired sequence. ∎
5 Impossibilities of $(\aleph_{0},2)$-Theorem for rectangles and triangles
In this section, we will prove theorem 1.17.
See 1.17

### v5

@@ 3.2 Monochromatic $(\aleph_{0},k+2)$-theorem implies the heterochromatic version
See 1.19
@@
Next, we shall show the tightness of the above result, that is Proposition 1.20. See 1.20
@@
Suppose $\mathcal{F}_{1}$ is the family containing only one unit ball $B_{1}$ in $\mathbb{R}^{2}$ with center at the origin and $\mathcal{F}_{2}$ contains the $4$ unit balls $B_{2,1},\dots,B_{2,4}$ in $\mathbb{R}^{2}$ with centers at the points $(1.5,0)$, $(-1.5,0)$, $(0,1.5)$, and $(0,-1.5)$ in $\mathbb{R}^{2}$ respectively. Then $\forall B\in\mathcal{F}_{2},\;B\cap B_{1}\neq\emptyset$ and clearly each of the families $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ are pierceable by a finite number of points. Now for each $n\in\mathbb{N},\;n\geq 3$, we define $\mathcal{F}_{n}$ to be the family of unit balls $\{B_{n,j}\;|\;j\in\mathbb{N}\}$, where $B_{n,j}$ is the unit ball in $\mathbb{R}^{d}$ with center at $c_{n,j}=(4j,4n),\;\forall j\in\mathbb{N}$. Then for any $m,n,i,j\in\mathbb{N},\;m,n\geq 3$ with $(m,i)\neq(n,j)$ we have $B_{m,i}\cap B_{n,j}=\emptyset$ and for each $n\in\mathbb{N},\;n\geq 3$, the family $\mathcal{F}_{n}$ is not pierceable by finite number of points. Hence $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ satisfies both the properties required. ∎
5 Impossibility results
In this section, we will prove Theorem 1.15, Theorem 1.16, Theorem 1.18 and Theorem 1.21.
See 1.15
@@ 5 Impossibility results
The construction here is similar to the construction in [KP22]. First we pick a ball $B_{1}=B(c_{1},r_{1})$. Next we pick another ball $F_{2}=B(c_{2},r_{2})$ such that $B_{1}\subset B(c_{2},\rho r_{2})$ and $B_{1}\cap F_{2}=\emptyset$. Let ¯$\bar{\ell}_{2}$ be a compact curve joining $c_{2}$ and any point of $B_{1}$. Define $B_{2}=F_{2}\cup\bar{\ell}_{2}$. In general, after $B_{1},B_{2},\dots,B_{m}$ are chosen, we pick another ball $F_{m+1}=B(c_{m+1},r_{m+1})$ such that $\cup_{i\in[m]}B_{i}\subset B(c_{m+1},\rho r_{m+1})$ and $\cup_{i\in[m]}B_{i}\cap F_{m+1}=\emptyset$. Let ¯$\bar{\ell}_{m+1}$ be a compact curve from $c_{m+1}$ that passes through $B_{i}$ for every $i\in[m]$ and does not intersect $B_{i}\cap B_{j}$ for any $i,j\in[m]$, $i\neq j$. Define $B_{m+1}=F_{m+1}\cup\bar{\ell}_{m+1}$ (see Figure 2). Then clearly, $\forall i\in[m],\;B_{m+1}\cap B_{i}\neq\emptyset$, i.e, $\{B_{n}\}_{n\in\mathbb{N}}$ satisfies the $(\aleph_{0},2)$ property. Again $\forall i,j\in[m]\text{ with }i\neq j,\;B_{m+1}\cap B_{i}\cap B_{j}=\emptyset$, i.e, no $3$ members of $\{B_{n}\}_{n\in\mathbb{N}}$ have a common point. So $\{B_{n}\}_{n\in\mathbb{N}}$ is not pierceable by a finite number of points. ∎
Figure 2: Sequence of non-convex sets with bounded condition number
@@
We will now prove Theorem 1.18.
Figure 6: One-unbounded family satisfying $(\aleph_{0},3)$-property with respect to lines without any finite line transversal.
See 1.18
Proof.
Recall the set of rectangles $\mathcal{F}$ from Theorem 1.16, where =$\mathcal{F}=\{R_{n}\}_{n\in\mathbb{N}}$. Now, for each $n\in\mathbb{N}$, lift $R_{n}$ to the plane $z=n$. Call the set of these lifted rectangles $\mathcal{F}^{\prime}$. Clearly, $\mathcal{F}^{\prime}$ does not have a finite hitting set consisting of lines. Now we show that every infinite subset of $\mathcal{F}^{\prime}$ has $3$ sets that are pierced by a single straight line. Let $\mathcal{S}\subset\mathcal{F}^{\prime}$ be an infinite subset of $\mathcal{F}^{\prime}$. For every $R\in\mathcal{S}$, choose a point $p_{R}\in R$. Let $q_{R}$ be the projection of $p_{R}$ onto the $xy$-plane. The set $\{q_{R}\}_{R\in\mathcal{S}}$ has a limit point $q$. Let $\vec{r}$ be a ray parallel to the $z$-axis with origin at $q$, where $\vec{r}$ is obtained by shifting the positive $z$-axis to $q$. For every $n\in\mathbb{N}$, $Q(\vec{r},n)$ (as defined in Section 2) contains $p_{R}$ for infinitely many $R\in\mathcal{S}$. Choose any two $R_{i},R_{j}\in\mathcal{S}$ and project them onto the $xy$-plane, and pick a point $q^{\prime}$ that is in the interior of the intersection of the two projected sets. Let $\vec{r}^{\prime}$ denote the ray from $q^{\prime}$ that is parallel to and is in the direction of $vec{r}$. Clearly, there is a cone $C$ with its endpoint at $q$ and $\vec{r}^{\prime}$ as its axis such that every ray in $C$ originating from $q^{\prime}$ intersects both $R_{i}$ and $R_{j}$ (see Figure 6). Since there is an $m\in\mathbb{N}$ for which $C\subset Q(\vec{r},m)$, the result follows. ∎
And we conclude this section by proving Theorem 1.21.
See 1.21
Figure 7: Dotted region denotes $A_{\{j_{1},j_{2}\}}$
Proof.